the perimeter of rectangle is 120ft. the length of rectangle is twice the width. find the length and the width of the rectangle.
the perimeter of traingle is 110 cm. one side is twice the second side. the third side is 30 cm more the than second side. find the length of each side.

Answer by mananth(16949)

let width be x
length of rectangle = 2x
,,
Perimeter of rectangle = 2L+2W
2x+4x=120
6x=120
x=20 feet which is the width
..
length = 2x = 2*20 =40 feet
..
let second side be x
first side will be 2x
Third side = x+30
Perimeter = 110
..
Perimeter = aum of all sides
x+2x+x+30= 110
4x+30 =110
4x=80
x=20 cm - second side
first side = 2x = 40cm
third side = x+30 = 20+30=50 cm